# A mechanism to generate varying speed of light via Higgs-dilaton coupling:Theory and cosmological applications

### v2

@@ Abstract
We consider a Higgs–dilaton action which allows the Higgs field $\Phi$ to interact with a dilaton field $\chi$ of the background spacetime through the term $\chi^{2}\,\Phi^{\dagger}\Phi$. Upon spontaneous gauge symmetry breaking, the vacuum expectation value (VEV) of the Higgs field becomes proportional to $\chi$. While traditionally this linkage is employed to make the Planck mass and particle masses dependent on $\chi$, we present an alternative mechanism: the Higgs VEV will be used to construct Planck’s quantum of action $\hbar$ and speed of light $c$. Specifically, each open set vicinity of a given point $x^{*}$ on the spacetime manifold is equipped with a replica of the Glashow–Weinberg–Salam action operating with its own effective values of $\hbar_{*}$ and $c_{*}$ per $\hbar_{*}\propto\chi^{-1/2}(x^{*})$ and $c_{*}\propto\chi^{1/2}(x^{*})$, causing these “fundamental constants” to vary alongside the dynamical field $\chi$. Moreover, in each open set around $x^{*}$, the prevailing value $\chi(x^{*})$ determines the length and time scales for physical processes occurring in this region as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. This leads to an anisotropic relation $\tau^{-1}\propto l^{-3/2}$ between the rate of clocks and the length of rods, resulting in a distinct set of novel physical phenomena. For late-time cosmology, the variation of $c$ along the trajectory of light waves from distant supernovae towards the Earth-based observer necessitates modifications to the Lemaître redshift relation and the Hubble law. These modifications are capable of: (1) Accounting for the Pantheon Catalog of Type Iasupernovae through a declining speed of light in an expanding Einstein–de Sitter universe, thus avoiding the need for dark energy; (2) Revitalizing Blanchard–Douspis–Rowan-Robinson–Sarkar’s CMB power spectrum analysis that bypassed dark energy [A&A 412, 35 (2003)]; and (3) Resolving the $H_{0}$ tension without requiring a dynamical dark energy component.
@@ I Motivation
In light of several outstanding issues and tensions in cosmology, growing interest has emerged in promoting the parameters of physical models to be variable scalar fields in spacetime [1, 3, 2]. For example, the cosmological constant $\Lambda$ is equipped with a kinetic term that allows it to evolve and ‘relax’ to its (small) value in the current epoch, thus potentially offering a relief to the fine-tuning and coincidence problems in late-time cosmology [5, 4].
Against this backdrop, the accustomed Planck $cGh$ unit system (1899-1900) solidifies the role of the speed light $c$, the (Newton) gravitational constant $G$, and the quantum of action $\hbar$ as “fundamental constants” which encompass the realms of special relativity, gravitation, and quantum mechanics, respectively [6, 7]. The trio $\{c,\,G,\,\hbar\}$ are viewed as fixed cornerstones, convertible to the three “fundamental units”—the Planck mass, the Planck length, and the Planck time, defined as
$M_{P}:=\sqrt{\frac{\hbar c}{G}};\ \ l_{P}:=\sqrt{\frac{\hbar G}{c^{3}}};\ \ \tau_{P}:=\sqrt{\frac{\hbar G}{c^{5}}}\,.$ (1)
@@
In the presence of Higgs–dilaton coupling, when the electroweak gauge symmetry of the action undergoes spontaneous breaking, the field $\chi$ influences the vacuum expectation value (VEV) of the Higgs doublet $\Phi$, which in turn affects the mass of the fermions, the gauge vector bosons, and the Higgs bosons in the GWS model. Specifically, Fujii and Wetterich allowed the particle mass $m$ to scale proportional to $\chi$, thereby maintaining a constant ratio $m/M_{P}$ at classical level. We shall refer to their practice collectively as the Fujii–Wetterich (FW) scheme. Moreover, the length scale $l$ and the time scale $\tau$ of any given physical process also become dependent on $\chi$ as $l\propto\chi^{-1}$ and $\tau\propto\chi^{-1}$, in exact proportion to $l_{P}$ and $\tau_{P}$ respectively. Importantly, the field $\chi$ affects not only the Planck time per Eq. (5) but also the rate of physical clocks (viz. the revolution rate of a mechanical clock, the oscillation rate of an atomic clock, or the decay rate of an unstable quantum system) which are made of matter governed by the GWS model. Quantitatively, the FW scheme thus predicts that the evolution rate $\tau^{-1}$ of physical processes is proportional to $\chi$.  [^2] We find it illuminating to quote Fujii [18]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [14]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
@@
Prior to the development of Brans-Dicke theory, Dicke also briefly considered variability in the speed of light $c$ in 1957 [20] although he did not pursue it further. A relatively little-known fact is that the concept of variable speed of light (VSL) was initially explored by Einstein in 1911, published in three obscure papers (originally in German) [21, 22, 23]  [^3] It seems that Dicke in 1957 [20] was not aware of Einstein’s 1911 VSL proposal [21]. during his quest for a generally covariant theory of gravity. In [22, 23] Einstein emphasized that the Michelson–Morley experimental results and Lorentz symmetry are meant to hold only locally; while $c$ is invariant with respect to local boosts, its value needs not be universal. Einstein envisioned the possibility that the gravitational potential affected the clock rate, hence generating a varying $c$ in spacetime. Such a variation in $c$ could then produce a refraction effect on light rays around the Sun’s disc as a result of Huygens’s principle [21]. The later successes of General Relativity (GR) quickly overshadowed Einstein’s 1911 paper on VSL, leading him to neglect the idea, however. In the 1990s, the VSL concept was revived independently by Moffat and by Albrecht and Magueijo to tackle issues in early-time cosmology, such as the horizon paradox [24, 25]. Since then, several researchers actively explore various aspects of VSL [53, 54, 29, 26, 47, 50, 49, 48, 37, 38, 67, 51, 52, 65, 66, 81, 95, 85, 63, 62, 61, 79, 55, 30, 76, 32, 87, 91, 69, 88, 64, 45, 34, 84, 92, 89, 93, 73, 80, 82, 86, 57, 56, 70, 90, 94, 58, 71, 68, 72, 96, 46, 44, 41, 83, 78, 33, 40, 39, 36, 35, 75, 74, 42, 60, 59, 77, 97, 98, 99, 100, 43, 31, 27, 28].
@@ II Higgs–dilaton coupling
$\mathcal{L}_{\text{grav}}=\chi^{2}\,\mathcal{R}-4\omega\,\nabla^{\mu}\chi\nabla_{\mu}\chi-V(\chi)\,.$ (12)
@@ III A new mechanism to generate variable $\hbar$ and $c$
$\int d^{4}x\,\biggl{[}i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}A_{\mu}\psi+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\biggr{]}\,.$ (14)
@@
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}\psi\,.$ (16)
@@
$\biggl{(}i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar c}}\,\gamma^{\mu}A_{\mu}+m\,\frac{c}{\hbar}\biggr{)}\,\psi=0\,.$ (17)
@@
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\frac{e}{\sqrt{\hbar c}}\,\bar{\psi}\gamma^{\mu}\psi\,.$ (18)
@@
$m:=\frac{f\mu}{\sqrt{\lambda}};\ \ \ e:=\sqrt{\alpha}\,.$ (19)
@@
$\mathcal{L}_{\text{QED}}=i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\frac{e}{\sqrt{\hbar_{*}c_{*}}}\bar{\psi}\gamma^{\mu}A_{\mu}\psi+m\frac{c_{*}}{\hbar_{*}}\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\,.$ (22)
@@
$\hat{H}(t)$ $:=-i\alpha^{k}\frac{\partial}{\partial x^{k}}+\sqrt{\alpha}\left(\alpha^{k}A_{k}-\beta A_{0}\right)+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\beta\,.$ (24)
@@ IV A prediction: Anisotropic scaling in the clock rate
$dx^{\mu}\rightarrow\chi_{*}\,dx^{\mu};\ \psi\rightarrow\chi_{*}^{-3/2}\psi;\ A_{\mu}\rightarrow\chi_{*}^{-1}A_{\mu}\,.$ (25)
@@
$E_{n=2}^{j=3/2}$ $=\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,.$ (32)
@@
$=\frac{f\mu}{2\pi\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}^{3/2}\,.$ (34)
@@
This predicted time dilation effect differs from the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric, arises from the dependence of the clock rate on the dilaton field $\chi$, viz. Eq. (36). This new phenomenon is distinct and shall be referred to as a “time dilation effect of the Third kind”, to be investigated in future work.  [^8] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer our predicted time dilation effect as “the Third kind”.
@@
$A=\frac{4\alpha}{3}\frac{m^{3}c_{*}^{4}}{h_{*}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}\,.$ (39)
@@
$A\propto\frac{\chi_{*}^{2}}{\chi_{*}^{-3/2}}\,\chi_{*}^{-2}\propto\chi_{*}^{3/2}\,.$ (40)
@@ V Comparison of our VSL scheme with the Fujii–Wetterich scheme
where “FW” refers to Fujii–Wetterich. This contrasts with our prediction, Eq. (27)
$\tau^{-1}\propto\chi_{*}^{3/2}\,.$
@@
$m_{\text{FW}}:=\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*};\ \ \ e:=\sqrt{\alpha}\,.$ (43)
@@
$\tau_{\text{FW}}\simeq\frac{\hbar_{\text{FW}}}{\hat{H}}\propto\chi_{*}^{-1}\,.$ (45)
@@ VI.1 VSL cosmology and late-time cosmography
    $F(z)=1+(1-F_{\infty})\bigl{(}1-(1+z)^{-2}\bigr{)}^{2}\,.$ (55)
@@
$z\approx\frac{3}{2}H_{0}\,\frac{d_{L}}{c_{0}}\,.$ (57)
@@
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}\,F(z)}\,\ln\frac{(1+z)^{2/3}}{F(z)}\,.$ (58)
@@
$\frac{d_{L}}{c}=\frac{1+z}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{M}(1+z^{\prime})^{3}+\Omega_{\Lambda}}}\,,$ (59)
@@
$\frac{d_{L}}{c}=2\,\frac{1+z}{H_{0}}\,\biggl{(}1-\frac{1}{\sqrt{1+z}}\biggr{)}\,.$ (60)

### v3

@@ Abstract
We probe into a class of scale-invariant actions, which allow the Higgs field $\Phi$ to interact with a dilaton field $\chi$ of the background spacetime through the term $\chi^{2}\,\Phi^{\dagger}\Phi$. Upon spontaneous gauge symmetry breaking, the vacuum expectation value (VEV) of the Higgs field becomes proportional to $\chi$. Although this linkage is traditionally employed to make the Planck mass and particle masses dependent on $\chi$, we present an alternative mechanism: the Higgs VEV will be used to construct Planck’s quantum of action $\hbar$ and speed of light $c$. Specifically, each open set vicinity of a given point $x^{*}$ on the spacetime manifold is equipped with a replica of the Glashow–Weinberg–Salam action operating with its own effective values of $\hbar_{*}$ and $c_{*}$ per $\hbar_{*}\propto\chi^{-1/2}(x^{*})$ and $c_{*}\propto\chi^{1/2}(x^{*})$, causing these “fundamental constants” to vary alongside the dynamical field $\chi$. Moreover, in each open set around $x^{*}$, the prevailing value $\chi(x^{*})$ determines the length and time scales for physical processes occurring in this region as $l\propto\chi^{-1}(x^{*})$ and $\tau\propto\chi^{-3/2}(x^{*})$. This leads to an anisotropic relation $\tau^{-1}\propto l^{-3/2}$ between the rate of clocks and the length of rods, resulting in a distinct set of novel physical phenomena. For late-time cosmology, the variation of $c$ along the trajectory of light waves from distant supernovae towards the Earth-based observer necessitates modifications to the Lemaître redshift formula, the Hubble law, and the luminosity distance–redshift relation. These modifications are capable of: (1) Accounting for the Pantheon Catalog of Type Ia supernovae through a declining speed of light in an expanding Einstein–de Sitter universe, thus avoiding the need for dark energy; (2) Revitalizing Blanchard–Douspis–Rowan-Robinson–Sarkar’s CMB power spectrum analysis that bypassed dark energy [A&A 412, 35 (2003)]; and (3) Resolving the $H_{0}$ tension without requiring a dynamical dark energy component.
@@ I Motivation
In light of several outstanding issues and tensions in cosmology, growing interest has emerged in promoting the parameters of physical models to be variable scalar fields in spacetime [1, 4, 2, 3]. For example, the cosmological constant $\Lambda$ is equipped with a kinetic term that allows it to evolve and ‘relax’ to its (small) value in the current epoch, thus potentially offering a relief to the fine-tuning and coincidence problems in late-time cosmology [5, 6].
Against this backdrop, the accustomed Planck $cGh$ unit system (1899-1900) solidifies the role of the speed light $c$, the (Newton) gravitational constant $G$, and the quantum of action $\hbar$ as “fundamental constants” which encompass the realms of special relativity, gravitation, and quantum mechanics, respectively [7, 8, 9]. The trio $\{c,\,G,\,\hbar\}$ are viewed as fixed cornerstones, convertible to the three “fundamental units”—the Planck mass, the Planck length, and the Planck time, defined as
$M_{P}:=\sqrt{\frac{\hbar c}{G}};\ \ l_{P}:=\sqrt{\frac{\hbar G}{c^{3}}};\ \ \tau_{P}:=\sqrt{\frac{\hbar G}{c^{5}}}$ (1)
@@
In the presence of Higgs–dilaton coupling, when the electroweak gauge symmetry of the action undergoes spontaneous breaking, the field $\chi$ influences the vacuum expectation value (VEV) of the Higgs doublet $\Phi$, which in turn affects the mass of the fermions, the gauge vector bosons, and the Higgs bosons in the Glashow–Weinberg–Salam (GWS) model. Specifically, Fujii and Wetterich allowed the particle mass $m$ to scale proportional to $\chi$, thereby maintaining a constant ratio $m/M_{P}$ at classical level. We shall refer to their practice collectively as the Fujii–Wetterich (FW) scheme. Moreover, the length scale $l$ and the time scale $\tau$ of any given physical process also become dependent on $\chi$ as $l\propto\chi^{-1}$ and $\tau\propto\chi^{-1}$, in exact proportion to $l_{P}$ and $\tau_{P}$ respectively. Importantly, the field $\chi$ affects not only the Planck time per Eq. (5) but also the rate of physical clocks (viz. the revolution rate of a mechanical clock, the oscillation rate of an atomic clock, or the decay rate of an unstable quantum system) which are made of matter governed by the GWS model. Quantitatively, the FW scheme thus predicts that the evolution rate $\tau^{-1}$ of physical processes is proportional to $\chi$.  [^2] We find it illuminating to quote Fujii [20]: “… the time and length in the microscopic unit frame are measured in units of $m^{-1}(t)$, in agreement with the physical situation that the time scale of atomic clocks, for example, is provided by the atomic levels which are determined by the Rydberg constant $(me^{4})^{-1}$ and Wetterich [16]: “The clock provided by the Hubble expansion in the standard description is now replaced by a clock associated to the increasing value of $\chi$.
@@
Prior to the development of Brans-Dicke theory, Dicke also briefly considered variability in the speed of light $c$ in 1957 [22] although he did not pursue it further. A relatively little-known fact is that the concept of variable speed of light (VSL) was initially explored by Einstein in 1911, published in three obscure papers (originally in German) [23, 24, 25]  [^3] It seems that Dicke in 1957 [22] was not aware of Einstein’s 1911 VSL proposal [23]. during his quest for a generally covariant theory of gravity. In [24, 25] Einstein emphasized that the Michelson–Morley experimental results and Lorentz symmetry are meant to hold only locally; while $c$ is invariant with respect to local boosts, its value needs not be universal. Einstein envisioned the possibility that the gravitational potential affected the clock rate, hence generating a varying $c$ in spacetime. Such a variation in $c$ could then produce a refraction effect on light rays around the Sun’s disc as a result of Huygens’s principle [23]. The later successes of General Relativity (GR) quickly overshadowed Einstein’s 1911 paper on VSL, leading him to neglect the idea, however. In the 1990s, the VSL concept was revived independently by Moffat and by Albrecht and Magueijo to tackle issues in early-time cosmology, such as the horizon paradox [26, 27]. Since then, several researchers actively explore various aspects of VSL [29, 30, 31, 28, 49, 52, 51, 50, 39, 40, 68, 53, 54, 66, 67, 82, 96, 86, 63, 62, 61, 80, 55, 32, 77, 34, 88, 92, 70, 89, 65, 47, 36, 85, 93, 90, 94, 74, 81, 83, 87, 57, 56, 71, 91, 95, 58, 72, 69, 73, 97, 48, 46, 43, 84, 79, 35, 42, 41, 38, 37, 76, 75, 44, 60, 59, 78, 98, 99, 100, 101, 45, 33, 64].
@@ II Higgs–dilaton coupling
$\mathcal{L}_{\text{grav}}=\chi^{2}\,\mathcal{R}-4\omega\,\nabla^{\mu}\chi\nabla_{\mu}\chi-V(\chi)$ (12)
@@ III A new mechanism to generate variable $\hbar$ and c
$\int d^{4}x\,\biggl{[}i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}A_{\mu}\psi+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\biggr{]}$ (14)
@@
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\sqrt{\alpha}\,\bar{\psi}\gamma^{\mu}\psi$ (16)
@@
$\biggl{(}i\gamma^{\mu}\partial_{\mu}+\frac{e}{\sqrt{\hbar c}}\,\gamma^{\mu}A_{\mu}+m\,\frac{c}{\hbar}\biggr{)}\,\psi=0$ (17)
@@
$\partial_{\nu}F^{\nu\mu}=j^{\mu}:=\frac{e}{\sqrt{\hbar c}}\,\bar{\psi}\gamma^{\mu}\psi$ (18)
@@
$m:=\frac{f\mu}{\sqrt{\lambda}};\ \ \ e:=\sqrt{\alpha}$ (19)
@@
$\mathcal{L}_{\text{QED}}=i\bar{\psi}\gamma^{\mu}\partial_{\mu}\psi+\frac{e}{\sqrt{\hbar_{*}c_{*}}}\bar{\psi}\gamma^{\mu}A_{\mu}\psi+m\frac{c_{*}}{\hbar_{*}}\bar{\psi}\psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ (22)
@@
$\hat{H}(t)$ $:=-i\alpha^{k}\frac{\partial}{\partial x^{k}}+\sqrt{\alpha}\left(\alpha^{k}A_{k}-\beta A_{0}\right)+\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}\,\beta$ (24)
@@ IV A prediction:   Anisotropic scaling in the clock rate
$dx^{\mu}\rightarrow\chi_{*}\,dx^{\mu};\ \psi\rightarrow\chi_{*}^{-3/2}\psi;\ A_{\mu}\rightarrow\chi_{*}^{-1}A_{\mu}$ (25)
@@
$E_{n=2}^{j=3/2}$ $=\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}\,\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*}$ (32)
@@
$=\frac{f\mu}{2\pi\sqrt{\lambda}}\Bigl{(}\sqrt{1-\text{\scriptsize$\frac{1}{4}$}\alpha^{2}}-\sqrt{1-\alpha^{2}}\Bigr{)}\,\chi_{*}^{3/2}$ (34)
@@
We emphasize that this predicted effect is physical, meaning it is, in principle, measurable by comparing the time lapses of two clocks situated at two separate locations with different values of the dilaton field. This time dilation effect differs from the time dilation effect in GR which is associated with the $g_{00}$ component of the spacetime metric, arises from the dependence of the clock rate on the dilaton field $\chi$, viz. Eq. (36). This new phenomenon is distinct and shall be referred to as a “time dilation effect of the Third kind”, to be investigated in future work.  [^8] In addition to the time dilation effect in GR, there is a well-known effect in Special Relativity where two twice-intersecting time-like paths can have different total amounts of proper time in between. Therefore, we refer our predicted time dilation effect as “the Third kind”.
@@
$A=\frac{4\alpha}{3}\frac{m^{3}c_{*}^{4}}{h_{*}^{3}}\left\langle\vec{r}_{if}\right\rangle^{2}$ (39)
@@
$A\propto\frac{\chi_{*}^{2}}{\chi_{*}^{-3/2}}\,\chi_{*}^{-2}\propto\chi_{*}^{3/2}$ (40)
@@ V Comparison of our VSL scheme with the Fujii–Wetterich scheme
where “FW” refers to Fujii–Wetterich. This contrasts with our prediction $\tau^{-1}\propto\chi_{*}^{3/2}$, given in Eq. (27).
@@
$m_{\text{FW}}:=\frac{f\mu}{\sqrt{\lambda}}\,\chi_{*};\ \ \ e:=\sqrt{\alpha}$ (43)
@@
$\tau_{\text{FW}}\simeq\frac{\hbar_{\text{FW}}}{\hat{H}}\propto\chi_{*}^{-1}$ (45)
@@ VI.1 VSL cosmology and late-time cosmography
    $F(z)=1+(1-F_{\infty})\bigl{(}1-(1+z)^{-2}\bigr{)}^{2}$ (55)
@@
$z\approx\frac{3}{2}H_{0}\,\frac{d_{L}}{c_{0}}$ (57)
@@
$\frac{d_{L}}{c_{0}}=\frac{1+z}{H_{0}\,F(z)}\,\ln\frac{(1+z)^{2/3}}{F(z)}$ (58)
@@
$\frac{d_{L}}{c}=\frac{1+z}{H_{0}}\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{M}(1+z^{\prime})^{3}+\Omega_{\Lambda}}}$ (59)
@@
$\frac{d_{L}}{c}=2\,\frac{1+z}{H_{0}}\,\biggl{(}1-\frac{1}{\sqrt{1+z}}\biggr{)}$ (60)